### v2

# Practical Secure Delegated Linear Algebra with Trapdoored Matrices

Abstract
Most heavy computation occurs on servers owned by a second party. This reduces data privacy, resulting in interest in data-oblivious computation, which typically severely degrades performance. Secure and fast delegated computation is particularly important for linear algebra, which comprises a large fraction of total computation and is best run on highly specialized hardware often only accessible through the cloud.
We state the natural efficiency and security desiderata for fast and data-oblivious delegated linear algebra. We demonstrate the existence of trapdoored matrix families based on a LPN assumption, and provide a scheme for secure delegated matrix-matrix and matrix-vectors multiplication based on the existence of trapdoored matrices. We achieve sublinear overhead for the server, dramatically reduced computation cost for the client, and various practical advantages over previous algorithms.
Keywords:
Data Privacy, Data-Oblivious Computation, Delegation, Homomorphic Encryption, Cloud Computing, Algorithm Efficiency, Sublinear Overhead, LPN, Matrix Multiplication, Trapdoored Matrix
1 Introduction
Data-oblivious delegated computation – computation in which the server does not learn anything meaningful about the input data, thanks to cryptographic scrambling by the client – has been of theoretical and practical interest for some time [PF79, Ost90]. Arbitrary computation is quite difficult in this setting, requiring Fully Homomorphic Encryption (FHE) [RAD+78, Gen09] and currently incurring several orders of magnitude of overhead. As a result, much work on oblivious computation focuses on either advancing schemes for oblivious computation with more limited adversaries [MLS+13, ZE15], on developing and improving the runtime of FHE [FV12], or on Partially Homomorphic Encryption (PHE) [ElG85, GM84], which focuses on encryption that commutes with some operations.
A second paradigm, with different uses, has also emerged. Secure Multiparty Computation (MPC) [Yao86, GMW87], in which several non-colluding adversaries attempt computation on a union of datasets private to each server, has allowed for a variety of recent developments in hidden computation protocols. Here, a combination of informational and cryptographic techniques often offer efficiency improvements over corresponding PHE problems [MZ17, AFL+16, Kel20, KVH+21].
Much work in FHE and MPC has focused on lattice techniques. Learning With Errors (LWE), a survey of which is available in [Reg10], is the most well-known cryptographic problem in this area. Broadly speaking, the problem asks a user to distinguish a sequence of random values from a sequence of low-magnitude-noised inner products of known vectors $a_{i}$ against some unknown vector $s$. Learning Parity with Noise (LPN), a variant problem, instead asks that the noise be sparse.
In recent decades, an increasing fraction of computational power has been dedicated to matrix-vector and matrix-matrix multiplications. Specialized hardware to perform these operations is expensive and power-hungry, and is therefore unusually centralized. We often wish to allow low-power devices (such as personal computing hardware) to do large matrix-vector multiplications (in evaluating artificial neural networks, for instance) or to allow companies to rent computing power from centralized providers. However, input data must often be kept private (e.g. medical records, personal photos). We therefore desire schemes for delegated matrix-vector and matrix-matrix multiplication that are both secure and require minimal computational overhead – when even insecure computation is quite expensive, a constant-factor overhead is extremely costly. Techniques using exclusively operations native to modern GPUs/TPUs are preferred, as they benefit from extensive hardware and software optimization.
1.1 Our Contributions
We wish to conduct remote computations of the form $Av_{i}$ for some prespecified matrix $A$ and online vectors $v_{i}$, or matrix-matrix multiplications $AB$, while hiding $A,B,v_{i}$ from the remote server and minimizing local and total computational load. We state the natural efficiency and security desiderata, propose two trapdoored matrix objects, and provide an algorithm for the above linear-algebraic problems from such objects satisfying the aforementioned desiderata.
We then provide an instantiation of such objects, and an associated algorithm, under a standard LPN [Kea98, HKL+12] hardness assumption. We also implement and benchmark our algorithm. It is lightweight and easy to implement, secure over any LPN-amenable ring or field, and uses only standard vector operations, rendering it highly amenable to compiler and hardware acceleration. As a result, it achieves both theoretically and practically small complexity blowup over computation in the clear while substantially reducing client computation.
1.2 Related Works
[Sot16] proposed a different version of trapdoored matrices, including a candidate construction. [^1] In this work, the definition of trapdoored matrices is not made exact, but weaker constructions (lacking pseudorandomness) are sufficient in their treated context, while stronger criteria are required in ours. We were unaware of this work in a previous version, and apologize for the omission.
Several recent papers have focused directly on multiplication of matrices when one or both are encrypted, such as [JKLS18, BCH+24, LZ22, GQH+24]. [LZ22, BCH+24] in particular achieve efficient computation and decoding in practical regimes, but require expensive encoding. Many of these works also use LWE variants and incur a small accuracy error as a result.
LPN-type problems were first discussed in [Kea98], and generalized to LPN in [HKL+12]. Best-known quantitative hardness result for LPN are discussed in [LWYY24].
Oblivious linear function evaluation (OLE) was, to the best of our knowledge, first specialized as a problem of particular interest by [DGN+17], who provided improvements to the polynomial evaluation case of [NP99] (while we do not consider MPC or OLE directly, the following work represents some of the more advanced uses of LPN to date).[BCGI18, BCG+19, BCG+20b] demonstrated the usefulness of LPN for secure linear computation in MPC. In particular, [BCG+19] presented a MPC algorithm that can be turned into a vector-secure matrix-vector multiplication algorithm, but focused on goals corresponding to the MPC setting. [ADI+17, AK23] consider low-overhead MPC linear-algebra, including an algorithm obtaining constant-factor overhead for secure two-party arithmetic computation with honest adversaries. [CKR+20] used alternate techniques to achieve matrix multiplication with linear communication in an all-but-one-adversarial MPC framework.
Secure matrix-vector multiplication is useful in various contexts, including evaluation of neural networks [MZ17, BEDK19, MWCB23].
In a concurrent and independent work [VZ25], Vaikuntanathan and Zamir suggest a recursive construction of trapdoored matrices very similar to that of our Subsection 7 and use it to accelerate algorithms that use random matrices. Their construction can be used to slightly improve our server cost of matrix-matrix (removing the term in $\delta$) while approximately doubling client cost.
A very recent follow-up work [BCH+25] has since provided alternate fast-matrix-vectors and Trapdoored-Matrix constructions using novel cryptographic assumptions, enabling significant storage and performance improvements in some domains. We expect further improved Trapdoored-Matrix constructions, with associated improvements in linear algebra delegation.
2 Fast Remote Computation
We work within the model of secure delegated computation problems. We assume two parties: a client with private data (e.g. two matrices, or a matrix and an online stream of vectors) and a server. The client wishes to learn some public function (e.g. matrix product) of their data while computationally hiding it from the server.
Since we consider clients that are much less computationally able than servers, we are interested in protocols which achieve the following properties:
  • The aggregate computation is very low (at most low-factor linear in, and ideally asymptotically equal to, the naive computational cost of the function).
  • The aggregate computation by the client is much less than that required to compute the function on its own (and ideally, nearly linear in input size).
  • The protocol is zero-knowledge for dishonest servers.
  • The client can detect dishonest server behavior.
  • The protocol minimizes communication and round count.
For security, we use the standard definition:
Definition 1
We say that a protocol is $(T(\cdot,\cdot),\epsilon(\cdot,\cdot),N(\cdot))$-secure if it has a security parameter $k$ such that, for all $n\geq N(k)$, the distributions of server views generated by any two client inputs are $(T(n,k),\epsilon(n,k))$-indistinguishable.
For efficiency, we compare to the unencrypted cost of the function:
Definition 2
We say that a protocol is $(c_{c}(n,k),c_{s}(n,k))$-remote-efficient if the client and the server expend $c_{c}(n,k)$ and $c_{s}(n,k)$ times the computation required to compute the function without encryption.
An ideal protocol would have $c_{c}(n,k)=0,c_{s}(n,k)=1$; an efficient protocol has $c_{c}(n,k)=o(1)$ and $c_{s}(n,k)=1+o(1)$
3 Three Problems and One Solution
We consider two particularly salient problems of linear algebra:
Definition 3 (Matrix-Matrix)
Given $A\in R^{m\times n},B\in R^{n\times l}$ over a ring $R$, return $AB$.
Definition 4 (Matrix-Vectors)
Given a $m\times n$ matrix $M$ over a ring $R$ and an online stream $v_{1},v_{2},\dots\in R^{n}$, return $Mv_{1},Mv_{2},\dots$ online.
Note: a secure remote protocol for Matrix-Vectors yields a natural protocol for Matrix-Matrix (by multiplying one column of $B$ at a time), but that protocol does not enjoy the same remote efficiency as the Matrix-Vectors protocol (as the unencrypted complexity of Matrix-Matrix is lower than $n$-repeated Matrix-Vectors).
These jointly generalize to the following third problem:
Definition 5 (Matrix-Matrices)
Given $A\in R^{m\times n}$ and an online stream of matrices $B_{i}\in R^{n\times l}$ over a ring $R$, return the online stream $AB_{i}$.
Our results depend on a LPN hardness assumption, and have performance directly proportional to the strength of that assumption. To state performance concisely, let $\nu(\delta,\epsilon,\lambda)=\max\left(\frac{\lambda}{\lg|R|},\nu\right)$, where $R$ is the ring over which we compute and $\nu$ is the smallest value such that LPN in dimension $\frac{\nu}{\delta}$ with subspace dimension $\nu$ and noise rate $\left(\frac{\nu}{\delta}\right)^{\epsilon-1}$ has $\lambda$-bit security.
We then show the following:
Theorem 3.1
For any ring $R$ over which the LPN assumption holds, Algorithm 2 is a secure delegated Matrix-Matrices algorithm with parameters $\delta,\epsilon$ restricted to domains where the LPN assumption holds and security parameter $\lambda$. It guarantees:
  • $\lambda$-bit-security CZK.
  • Preprocessing costs
    $O\left(n\left(n^{\epsilon}(m+n)+\nu(\delta,\epsilon,\lambda)(\delta n+m)\right)\right)$
    and
    $O\left(n^{\omega}\delta^{2\omega-3}+\min(m,\delta n)^{\omega-2}n\max(m,\delta n)\right)$
    to the client and server respectively. [^2] Where $\omega$ is the matrix multiplication constant.
  • Per-round cost
    $O((m+n)l\nu(\delta,\epsilon,\lambda)+(m+n)n^{\epsilon}l)$
    for the client and per-round overhead (i.e. excluding the cost of a single naive-sized $m\times n$-by-$n\times l$ multiplication)
    $O(n\max(\delta n,l)\min(\delta n,l)^{\omega-2})$
    for the server.
  • Two-round preprocessing and one-round online steps.
  • Communication (both in preprocessing and in online steps) of $1+O\left(\delta\right)$ times input size.
In particular, in parameter regimes where $m=\Omega(n)$, the algorithm is
$\left(n^{\epsilon}+\frac{\nu(\delta,\epsilon,\lambda)}{n},1+O\left(\delta^{\omega-2}\right)\right)$
remote-efficient in both Matrix-Matrix and $\max(n^{\epsilon},\delta n^{\omega-2})$-steps-amortized Matrix-Vectors.
This specializes:
Corollary 1
Algorithm 2 restricts to a two-round [^3] A minor modification is required to parallelize independent server calls. secure delegated $n$-by-$n$ matrix-matrix protocol with parameter $\delta$ and client cost $O(n^{2}(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)))$ and server cost beyond a single $n$-by-$n$ matrix multiplication of $O\left(\delta n^{\omega}\right)$.
Corollary 2
Algorithm 2 restricts to a two-preprocessing-round, one-round-per-online secure delegated $n$-by-$n$ matrix-vectors protocol with parameter $\delta$ and:
  • Client online cost $O(n(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)))$
  • Client preprocessing cost $O(n^{2}(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)))$
  • Server online cost beyond a single $n$-by-$n$ matrix-vector multiplication of $O\left(\delta n^{2}\right)$
  • Server initialization cost of $O(\delta n^{\omega})$
4 Primitives for Fast Linear Algebra
Our approach to security is a delegation-oriented modification of Beaver’s triples [Bea92]: to secure the computation $AB_{i}$, generate pseudorandom matrices $A^{\prime}$ and $B^{\prime}_{i}$. Request the computation $(A+A^{\prime})(B_{i}+B^{\prime}_{i})$ from the server, and return $(A+A^{\prime})(B_{i}+B^{\prime}_{i})-A^{\prime}B_{i}-(A+A^{\prime})B^{\prime}_{i}$ (where the latter two terms are computed locally). While this is clearly secure by the same argument, it is not clearly efficient: it requires two local matrix multiplications. To improve efficiency, we must pick $A^{\prime}$ and $B^{\prime}_{i}$ to reduce the cost of these operations, inspiring the following cryptographic primitive generalizing an idea of [Sot16].
Definition 6 (Trapdoored-Matrix)
A $(T(\cdot,\cdot),\epsilon(\cdot,\cdot),N(\cdot))$-secure, $C(\cdot,\cdot,\cdot)$-efficient Trapdoored-Matrix scheme is an algorithm that, given a ring $R$ and $m,n,l,k\geq N((m,n,R))\in\mathbb{Z}$, generates a $(T((m,n,R),k),\epsilon((m,n,R),k))$-pseudorandom $m\times n$ matrix $M$ over $R$ and $\tilde{O}(C(m,n))$ additional data such that, for any $N\in R^{n\times l}$, $MN$ can be computed in time $C(m,n,l)$. [^4] Note that this assumption is stronger than the fast-pseudorandom-matrix assumption of [ADI+17]: that assumption only requires that images under the matrix are random once noised. Many constructions satisfying that assumption (e.g. any sparse matrix) are insufficient here.555[Sot16] proposes a construction under a nonstandard security assumption. A construction of [BCG+20b] (sec. 10.3; the first construction with the Toeplitz/quasi-cyclic assumption) also fulfills these criteria, albeit with significantly stronger LPN assumptions and worse performance at equal security levels (as it was optimized for a substantially different task).
This scheme suffices for matrix-matrix multiplication. However, we wish also to hide matrix-vector multiplication, and therefore to develop pseudorandom trapdoored-vectors whose product with an arbitrary matrix $A$ can be computed in time much less than the naive product. This is impossible for one-shot instances, as the time to compute $Av$ for arbitrary $v$ is asymptotically equal to the time required to read $A$; therefore, we consider a further generalization of the above.
Definition 7 (Targeted-Trapdoored-Matrix)
A $(T(\cdot,\cdot),\epsilon(\cdot,\cdot),N(\cdot))$-secure, $C(\cdot,\cdot,\cdot)$-efficient Targeted-Trapdoored-Matrix scheme with $D(\cdot,\cdot,\cdot)$ preprocessing is an algorithm that, given a ring $R$, $m,n,l,k\geq N((m,n,R))\in\mathbb{Z}$, and a matrix $B\in R^{n\times l}$, preprocesses $B$ in time $D(\cdot,\cdot,\cdot)$, returning a $\tilde{O}(D(m,n))$-sized generator that generates $(T((m,n,R),k),\epsilon((m,n,R),k))$-pseudorandom $m\times n$ matrices $M_{i}$ and the associated value $M_{i}B$ in time $C(m,n,l)$.
Combined, these make the template scheme for Matrix-Matrix fast even when $B_{i}$ is a vector: use the Trapdoored-Matrix generator to generate $A^{\prime}$, and use the Targeted-Trapdoored-Matrix generator to generate the various $B^{\prime}_{i}$.
Assuming a large cache, we can obtain amortized Targeted-Trapdoored-Matrix from Trapdoored-Matrix: generate a trapdoored matrix $M$, compute $AM$, and then use blocks of $M$ as the $B^{\prime}_{i}$. Since we already need to store $A$, this space overhead is constant-factor.
We know of many families of matrices (Vandermonde matrices, Cauchy Matrices, discrete Chebyshev matrices) for which multiplication is $\tilde{O}(n)$. Since linear combinations of these or other fast-multiplication matrices could plausibly be pseudorandom, it seems natural to believe the following:
Conjecture 1
There exist Trapdoored-Matrix constructions with preprocessing cost $\tilde{O}(n^{2})$ and matrix-vector multiplication cost $\tilde{O}(n)$.
The complexity of Targeted-Trapdoored-Matrix in the case where $m=1$ (i.e. Targeted-Trapdoored-Vectors) is less clear. The security assumption of Section 5 is commonly believed to hold for any $\delta,\epsilon>0$. Security of LPN in the dimension regime required for polylogarithmic overhead, however, is less plausible; if a polylogarithmic-overhead scheme exists, it will likely require another approach.
5 A LPN Hardness Assumption and Performance Scaling
Given a ring $R$, a dimension $n$, a subspace dimension $n_{1}$, and a noise rate $\mu$, the decisional LPN problem asks one to distinguish messages of the form $(L,Lr+s)$ from messages of the form $(L,u)$ where $L\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n\times n_{1}}$, $r\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{1}}$, $u\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n}$, and $s\leftarrow\mathcal{D}_{R,\mu}^{n}$ where $\mathcal{D}_{R,\mu}$ is the distribution that is 0 with probability $1-\mu$ and uniform over $R$ with probability $\mu$. A standard hardness assumption for LPN (as in [Ale03, BCGI18, YZ21]) is the following:
Conjecture 2 (Standard LPN Hardness Assumption)
$\exists\epsilon<1$ such that fixing any $\delta>0$, there exists $n^{*}\in\mathbb{N},c>0$ such that for any $n>n^{*},n_{1}=\delta n$, $\mu=n^{\epsilon-1}$, no algorithm running in time $T=2^{O_{\delta}(n^{c})}$ can solve decisional-LPN with parameters $n,n_{1},\mu$ with advantage $\geq 2^{-\Omega_{\delta}(n^{c})}$.
Both $\delta$ and $\epsilon$ directly correspond to the parameters by the same names in Theorem 3.1. In particular, slightly stronger assumptions directly translate into significant improvements to the efficiency of the algorithm. For instance, if $\delta$ is allowed to be $o(1)$, we achieve sub-multiplicative overhead for $\frac{m}{n}\geq\delta$, rather than just $m=\Omega(n)$. All terms of the form $n^{\epsilon}$ correspond to $n\mu$; if we assume that $\mu(n)=n^{\epsilon-1}$ is secure for any $\epsilon>0$, we obtain the corresponding substantial efficiency improvements. Likewise, the security of our protocol depends directly on the computational hardness assumption: a super-polynomial assumption would yield a protocol with super-polynomial security.
Note: our matrix-multiplication results are nontrivial when $\epsilon<\omega-2$. The above assumption is commonly believed to hold (nonuniformly) for all $\epsilon$, or at least for $\epsilon$ quite small.
6 A Simple Secure Delegated Matrix-Matrices Algorithm
We first use LPN to construct schemes for Trapdoored-Matrix and Targeted-Trapdoored-Matrix, and therefore Secure Delegated Matrix-Matrices. These, and an associated algorithm for Secure Delegated Matrix-Matrices, are detailed in Algorithm 1. Our approach is simple: we observe that a matrix of the form $HL+S$, where $H$ and $L$ have few columns and rows respectively and $S$ is sparse, is pseudorandom by the LPN assumption and is easy to multiply by other matrices. Moreover, given a matrix $A$, once $AL^{\top}$ has been computed, it is very efficient to compute $A(LH+S)$ as $(AL)H+AS$, even if $LH+S$ represents a thin matrix (or vector).
A brief note on notation: in order to aid the reader in tracking the values represented by intermediate variables, we use angled brackets around an expression, such as $\langle\mathrm{Expr}(x,y,z,\dots)\rangle$ to denote a variable whose value is equal to that of the expression $\mathrm{Expr}(x,y,z,\dots)$.
Algorithm 1 Basic Secure Delegated Matrix-Matrices via Trapdoored-Matrix and Targeted-Trapdoored-Matrix Schemes
Procedures are written for the client. The $\mathbf{Send}$ and $\mathbf{SMult}$ commands are used, respectively, to send information and to request matrix multiplications (sending their arguments if the server does not already have them) to/from the server. For efficiency, $L_{i}$s are shared over all routines.
1:procedure TM-Init
2:     Input: $m,n\in\mathbb{N},n_{1}\in\mathbb{Z}_{>0},\mu\in\mathbb{R}_{>0}$
3:     $L\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{1}\times n}$, $H\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{m\times n_{1}}$, $S\overset{{\scriptscriptstyle\$}}{\leftarrow}\mathcal{D}_{\mu}^{m\times n}$
4:     $A^{\prime}=HL+S$
5:     return $A^{\prime}$
6:procedure TM-Mult
7:     Input: $M\in R^{n\times l}$
8:     return $(SM)+(H(LM))$triangleright Equal to $A^{\prime}M$
9:procedure TTM-Init
10:     Input: $n\in\mathbb{N},n_{1}\in\mathbb{Z}_{>0},\mu\in\mathbb{R}_{>0},N\in R^{m\times n}$
11:     $L_{R}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{1}\times n}$
12:     $P=NL_{R}^{\top}$
13:procedure TTM-Gen-Mult
14:     Input: $l\in\mathbb{N}$
15:     $H_{R}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{l\times n_{1}}$, $S_{R}\overset{{\scriptscriptstyle\$}}{\leftarrow}\mathcal{D}_{\mu}^{l\times n}$
16:     return $(L_{R}^{T}H_{R}^{T}+S_{R}^{T},PH_{R}^{T}+NS_{R}^{T})$$\triangleright$ Equal to $B^{\prime}_{i}$ and $NB^{\prime}_{i}$, respectively
17:procedure HFMM-Init
18:     Input: $A\in R^{m\times n}$, $n_{1}\in\mathbb{Z}_{>0}$, $\mu\in\mathbb{R}_{>0}$
19:     $A^{\prime}=\textsc{TM-Init}(m,n,n_{1},\mu)$
[…]
21:procedure HFMM-Mult
22:     Input: $B_{i}\in R^{m\times l}$
23:     $\langle A^{\prime}B_{i}\rangle=\textsc{TM-Mult}(B_{i})$
24:     $(B^{\prime}_{i},\langle(A+A^{\prime})B^{\prime}_{i}\rangle)=\textsc{TTM-Gen-Mult}(l)$
25:     return $\mathbf{SMult}(\langle A+A^{\prime}\rangle,\langle B_{i}+B^{\prime}_{i}\rangle)-\langle A^{\prime}B_{i}\rangle-\langle(A+A^{\prime})B^{\prime}_{i}\rangle$
6.1 Security and Performance
The security of this scheme is a direct consequence of the LPN assumption: the matrices $A^{\prime}$ and $B^{\prime}$ that TM-Mult and TTM-Mult generate have rows or columns constructed as LPN pseudorandom vectors, and therefore $A+A^{\prime},B+B^{\prime}$ (the only information that the server receives) leak no information about $A$ and $B$.
[…]
For most parameter regimes (e.g. $\delta=\Theta(1),\mu=o(1)$), the dominant cost comes from the dense multiplication $HL$, and therefore varies polynomially in subspace dimension $n_{1}$. This can be greatly reduced via LPN variants over fast matrices (e.g. sparse-LPN, Toeplitz-LPN, Quasi-cyclic-LPN [AMBD+18]), but parameter tradeoffs for these problems over large rings are not currently well-understood, and may entail worse $\mu$. In the next subsection, we present an alternate option that obtains equal or better performance (depending on the comparative sparsity assumptions of LPN and sparse-LPN) than a sparse-LPN version, but depends only on standard LPN assumptions.
In the case where $A$ is public, we can also achieve substantial improvements via dual-LPN. This simple substitution (i.e. calculating $v^{\prime}=Ps$, $Av^{\prime}=(AP)s$ for a public random square matrix $P$) takes the cost down to $O(mn\mu)$. The dual formulation will not, however, be sufficient to achieve very fast fully private matrix-matrices.
7 An Improved Secure Delegated Matrix-Matrices Algorithm
The efficiency of the previous section’s algorithm is primarily determined by the $\delta$ of the LPN assumption, which is rather high. This is not (just) an artifact of our weak assumption: the security of LPN does not degrade gracefully as its parameters are reduced. In particular, once $n_{1}\mu\leq 1$, straightforward attacks from subsampling and Gaussian elimination are feasible. This increases the overhead of Algorithm 2 for large inputs. In this section, we present a more complicated algorithm that bypasses this problem to improve performance.
The client cost of our earlier scheme is dominated by the costs of the multiplications $AL$ and $(AL)r_{i}$. There is a natural solution: recursively call the remote server to compute these while not leaking information about $A$ or $r_{i}$. Additionally, computations where both matrices are public (in particular, those of the form $L(B+B^{\prime})$ and similar) may be offloaded to the server. These changes result in a fairly complicated recursive protocol with several disadvantages, including a number of rounds growing exponentially in the recursion depth and substantial duplication of work, but this protocol in turn may be greatly optimized and simplified. We detail the resulting algorithm for secure delegated matrix-matrices in Algorithm 2, and spend the rest of the section demonstrating that it satisfies Theorem 3.1.
Algorithm 2 Client Routines for LPN-based Matrix-Vectors
1:procedure Mat-Mul-Check
2:     Input: $A\in R^{a\times b},B\in R^{b\times c},C\in R^{a\times c}$, $\lambda^{\prime}\in\mathbb{N}$ $W\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{c\times\lambda^{\prime}}$
3:     if $A(BW)\neq CW$ then Terminate
4:procedure Get-Checked-Cum-Prods
5:     Input: $M_{i}\in R^{a_{i}\times a_{i+1}}\forall i\in[k]$
6:     for $i\in\{2,3,\dots,k\}$ do
7:         $\langle M_{1}M_{2}\dots M_{i}\rangle=\mathbf{SMult}(\langle M_{1}M_{2}\dots M_{i-1}\rangle,M_{i})$
[…]
9:procedure Right-Hidden-Init
10:     Input: $M\in R^{s\times n}$, $n_{i}\forall i\in[d],\mu_{i}\forall i\in[d]$
[…]
12:procedure Right-Hidden-Mult
13:     Input: $V\in R^{n\times o}$.
14:     $H\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{d}\times o}$
15:     $\forall i\in[d],S_{i}\overset{{\scriptscriptstyle\$}}{\leftarrow}\mathcal{D}_{\mu_{i}}^{n_{i-1}\times o}$
16:     $V^{\prime}\leftarrow\langle L_{1}L_{2}\dots L_{d}\rangle H+\sum_{i=1}^{d}\langle L_{1}L_{2}\dots L_{i-1}\rangle S_{i}$
17:     $\langle MV^{\prime}\rangle\leftarrow\langle ML_{1}L_{2}\dots L_{d}\rangle H+\sum_{i=1}^{d}\langle ML_{1}L_{2}\dots L_{i-1}\rangle S_{i}$
[…]
19:     return $\langle M(V+V^{\prime})\rangle-\langle MV^{\prime}\rangle$.
20:procedure Hidden-Init
21:     Input: $A\in R^{m\times n}$, $n_{i}\forall i\in[d],\mu_{i}\forall i\in[d-1]$
22:     for $i\in[d]$ do $L_{i}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{i}\times n_{i-1}}$, $S_{i}\leftarrow\mathcal{D}_{\mu_{i}}^{m\times n_{i-1}}$      $H\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{m\times n_{d}}$.
23:     for $i\in[d]$ do $\mathbf{Send}(L_{i})$
[…]
25:     $C\leftarrow[\langle L_{1}\rangle,\langle L_{1}L_{2}\rangle,\dots,\langle L_{1}L_{2}\dots L_{d}\rangle]$triangleright as a block matrix
[…]
27:     $[\langle AL_{1}\rangle,\langle AL_{1}L_{2}\rangle,\dots,\langle AL_{1}\dots L_{d}\rangle]\leftarrow\textsc{Right-Hidden-Mult}(C^{\top},A^{\top})^{\top}$triangleright Divide $C^{\top}$ into a row-vector of blocks for better efficiency
28:     $A^{\prime}=H\langle L_{1}L_{2}\dots L_{d}\rangle^{\top}+\sum_{i=0}^{d-1}S_{i+1}\langle L_{1}L_{2}\dots L_{i}\rangle^{\top}$
29:     $\mathbf{Send}(A+A^{\prime})$
30:procedure Hidden-Mult
31:     Input: $B\in R^{m\times l}$.
32:     $R\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{d}\times l}$
33:     for $i\in[d]$ do $S_{i}\leftarrow\mathcal{D}_{\mu_{i}}^{n_{i-1}\times l}$
34:     $B^{\prime}\leftarrow\langle L_{1}L_{2}\dots L_{d}\rangle R+\sum_{i=1}^{d}\langle L_{1}L_{2}\dots L_{i-1}\rangle S_{i}$
35:     $\langle AB^{\prime}\rangle\leftarrow\langle AL_{1}L_{2}\dots L_{d}\rangle R+\sum_{i=1}^{d}\langle AL_{1}L_{2}\dots L_{i-1}\rangle S_{i}$
36:     for $i\in[d]$ do $\mathbf{SMult}((L_{1}L_{2}\dots L_{i})^{\top},(B+B^{\prime}))$.
37:     $\langle A^{\prime}(B+B^{\prime})\rangle\leftarrow H\langle(L_{1}L_{2}\dots L_{d})^{\top}(B+B^{\prime})\rangle+\sum_{i=0}^{d-1}S_{i+1}\langle(L_{1}L_{2}\dots L_{i})^{\top}(B+B^{\prime})\rangle$
38:     return $\mathbf{SMult}(\langle A+A^{\prime}\rangle,\langle B+B^{\prime}\rangle)-\langle AV^{\prime}\rangle-\langle A^{\prime}(B+B^{\prime})\rangle$
7.1 Algorithm Structure
Algorithm 2 is divided into several subroutines:
  • Mat-Mul-Check: Given two matrices $A$, $B$, and a third matrix $C$, check that $AB=C$ with negligible statistical soundness error.
  • Get-Checked-Cum-Prods: Given a sequence of matrices $M_{i}\in R^{a_{i}\times a_{i+1}}$ for a nonincreasing sequence of positive integers $a_{i}$, query the server for the sequence of cumulative products $M_{1}M_{2},M_{1}M_{2}M_{3},\dots$, and check the answer up to negligible statistical soundness error.
  • Right-Hidden-Init: Initialize a simplified version of the protocol that can offload a matrix-matrix multiplication to the server while hiding one operand.
  • Right-Hidden-Mult: Apply the above protocol.
  • Hidden-Init: Initialize the full protocol.
  • Hidden-Mult: Run the initialized protocol on a target operand.
All pseudorandom matrices are defined via recursion on LPN; for instance, we observe that
$V^{\prime}$ $=L_{1}L_{2}\dots L_{d}H+\sum_{i=1}^{d}L_{1}L_{2}\dots L_{i-1}S_{i}$
$=(S_{1}+L_{1}(S_{2}+L_{2}(S_{3}+L_{3}(\dots(S_{d}+L_{d}H)))))$
That this is pseudorandom will follow from a hybrid argument on the LPN assumption: that a parenthetical is pseudorandom, combined with LPN hardness, shows that the parenthetical containing it is pseudorandom.
Recursing on LPN makes the required dense matrix-vector multiplication substantially smaller (at the cost of an additional sparse multiplication), and so improves efficiency. This mostly decouples instance size from the security/cost tradeoff. In essence, we have a LPN variant whose security and computational cost both degrade gracefully in the subspace dimension.
7.2 Security and Correctness Guarantees
Our protocol makes three guarantees. First, in the semi-honest model, the protocol will return exactly $AB$ with probability 1. Second, the protocol guarantees computational zero-knowledge against even a dishonest server. Third, it guarantees a weaker form of correctness: that a server engaged in dishonesty in any setup query or fraction $\alpha$ of the online invocations can be detected with high probability with $O\left(\frac{1}{\alpha}\right)$ additional queries (and if $l$ and $|R|$ are large, much more efficient per-call checks are available via Mat-Mul-Check).
The first guarantee follows from expanding the algebra in the return line and noting that all products except $AB$ cancel. We emphasize that all LPN noise exactly cancels; unlike some LWE-based protocols, we introduce no error in the output.
For the security guarantees, we first notice that Mat-Mul-Check, a procedure first defined (to the best of our knowledge) in [Moh11], will return true if $AB=C$, but will return true with probability $|R|^{-\lambda^{\prime}}$ otherwise. Note that this provides low soundness error at very low computational cost for large rings (assuming that $b=O(a,c)$): for 32-bit integers, for instance, we achieve soundness error $2^{-128}$ when $\lambda^{\prime}=4$. Assuming that $\lambda^{\prime}$ is picked so that this probability is negligible, we demonstrate that Algorithm 2 is CZK.
Theorem 7.1
Given the LPN assumption of Section 5, Algorithm 2 is computational zero-knowledge for the server.
Proof
We construct a server-view simulator. The simulator is quite simple: it runs a copy of the client with inputs $A=0,B_{i}=0$. The $L_{i}$ generated by the simulator are identically distributed to those generated by the true client, as those are random. If the server reports any of the products $L_{1}L_{2}\dots L_{i}$ or $ML_{1}L_{2}\dots L_{i}$ falsely, both the simulator and a true client will detect the error and terminate the protocol with probability $\geq 1-|R|^{\lambda^{\prime}}$, so the complete views in this case are statistically indistinguishable. Otherwise, the rest of the client’s messages to the server consist exclusively of matrices taking the form of a hidden matrix $H$ plus an independently sampled matrix of the form $(L_{1}L_{2}\dots L_{d})R+\sum_{i=1}^{d}(L_{1}L_{2}\dots L_{i-1})S_{i}$ (where $R$ and $S_{i}$ are sampled as in Right-Hidden-Mult). The simulated client’s messages are identically composed, except that all of the matrices $H$ are zero. In other words, the total server view contains no information not fixed by $L_{1},\dots,L_{n}$ and a set of matrices decomposable into vectors generated as $h+(L_{1}L_{2}\dots L_{d})r+\sum_{i=1}^{d}(L_{1}L_{2}\dots L_{i-1})s_{i}$, where $h$ is 0 in the simulator and arbitrary but independent of the $L_{i}$s, $r$s, and $s_{i}$s in the true client (and $r$, $s_{i}$ are resampled for every such vector).
The following simple lemma shows that both such distributions are computationally indistinguishable from noise (and therefore from each other) conditioned on the LPN assumption, thus completing the proof.
Lemma 1
For $i\in[d]$, let $L_{i}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{i}\times n_{i-1}}$. From this, for $j\in\{0,1,\dots,d\}$, let
$\mathcal{E}_{j}=\{L_{1}L_{2}\dots L_{j}r_{j}+\sum_{i=1}^{j}(L_{1}L_{2}\dots L_{i-1})s_{i}|r_{j}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{j}},s_{i}\leftarrow\mathcal{D}^{n_{i-1}}_{\mu_{i}}\}.$
Say that, for all $j\in[i]$, the distributions
$\mathcal{R}_{i}=\left\{r_{i-1}\overset{{\scriptscriptstyle\$}}{\leftarrow}R_{n_{j-1}}\right\},\mathcal{L}_{i}=\left\{L_{i}r_{i}+s_{i}:r_{i}\overset{{\scriptscriptstyle\$}}{\leftarrow}R_{n_{i}},s_{i}\leftarrow\mathcal{D}^{n_{i-1}}\right\}$ (1)
are $(T,\epsilon_{j})$-indistinguishable conditioned on $L_{j}$ (as given by the LPN assumption). Then $\mathcal{E}_{0}$ and $\mathcal{E}_{d}$ are $(T-\mathrm{\,poly}(n,\mu),\sum_{i=1}^{d}\epsilon_{i})$-indistinguishable conditioned on $(L_{1},\dots,L_{j})$. [^6] Note that the offset term represents the cost of the conversion step in the distinguisher.
Proof
We prove the contrapositive via hybrid argument. Let there be an algorithm $\mathcal{A}$ distinguishing $\mathcal{E}_{0}$ and $\mathcal{E}_{d}$ with advantage $\geq\sum_{i=1}^{d}\epsilon_{i}$. Then there exists $j\in[d]$ s.t. $\mathcal{E}_{j-1}$ and $\mathcal{E}_{j}$ are distinguished by $\mathcal{A}$ with advantage $\geq\epsilon_{j}$. This implies an algorithm distinguishing the distributions $\mathcal{R}_{j}$ and $\mathcal{L}_{j}$ with advantage $\geq\epsilon_{j}$: sample matrices $L_{1},L_{2},\dots,L_{j-1}$ and noises $s_{1},\dots,s_{j-1}$ and apply them in the natural way to generate distributions $\mathcal{E}_{j-1}$ and $\mathcal{E}_{j}$, and then apply $\mathcal{A}$.
For the final security guarantee, we observe that the only place that the server can deviate from protocol without high-probability detection is in the last two queries (as all others are wrapped by Mat-Mul-Check. To check for $\alpha$-fraction of deviations in these queries, the client may introduce $\frac{c}{\alpha}$ multiplication queries with $B_{i}=0$. By linearity of the algorithm, the error introduced in either query propagates linearly into the result, and so if the server behaves dishonestly, the client will compute an incorrect result. The client may check this – $AB_{i}=0$ if $B_{i}=0$ – and given deviation rate $\alpha$, the probability of a deviation in $\frac{c}{\alpha}$ queries is $\geq 1-e^{-c}$, so any significant error rate can be detected cheaply and with very high probability.
7.3 Performance
See Appendix 0.A for asymptotic and per-operation performance exposition.
7.4 Practical Improvements
We proved bounds in the case where $\delta,\mu$ are constant across layers. This is inefficient; by shrinking $\delta,\mu$ in larger layers to get approximately uniform security, dramatic improvements (in the server’s $o(1)$ term and the client’s full cost) can be achieved.
Careful readers may note that this algorithm’s pseudorandom generation is very similar to a non-recursive instance of dual-LPN on the block syndrome matrix
$\left[L_{1},L_{1}L_{2},\dots,L_{1}L_{2}\dots L_{d}\right]$
It may initially appear that we can simply substitute dual-LPN for this entire process and gain major client performance and simplicity improvements. We can, but only at factor-two (over unencrypted computation) cost to the server: the ancillary product of line 27 becomes much larger. We can, however, achieve nearly the same improvements by substituting dual-LPN for the final recursive step (i.e. replacing the generation $L_{d}v+s_{d}$ by $Ps_{d}$). This entirely eliminates the client dense multiplication, which is quite costly in regimes where $n$ is small. This does increase the server overhead, but this increase is on the multiplicative order of $\delta^{d-1}$. As a result, client performance improves substantially (and server performance decreases marginally) when $n$ is close to $n_{d}$, which occurs when $n$ is low or particularly high security is desired.
8 Experiments
We implemented Algorithm 2 in C++ and benchmarked it for both matrix-matrix multiplication and repeated matrix-vector multiplication. All experiments were run on a single core of an AMD Ryzen 5700X3D processor. All parameters were chosen for 128-bit security, using updated versions of the estimates of [LWYY24]. Results are in Tables 1 and 2. Client R. and Total R. indicate the ratios of client and total compute time, respectively, to the unencrypted compute time (denoted Local). We noticed substantially worse performance when $n$ was set to a power of two (presumably due to cache bottlenecking); therefore, we set $n$ to be one more than a power of two. Benchmarks with $n$ exactly a power of two are available in Tables 3 and 4 of the appendix; global performance is substantially worse, but relative performance is consistent.
$n$ Local (s) Client Pre. (s) Server Pre. (s) Client (s) Server (s) Client R. Total R.
1025 0.00113 0.0000729 0.0000371 0.00117 0.00121 1.09 2.19
2049 0.00520 0.000145 0.0000830 0.00306 0.00527 0.616 1.65
4097 0.0570 0.000415 0.000388 0.00605 0.0576 0.113 1.13
8193 0.268 0.000927 0.00173 0.0145 0.270 0.0573 1.07
16385 1.27 0.00221 0.00467 0.0475 1.27 0.0392 1.04
Figure 1: Matrix-Vectors performance benchmarks. Client and server times separated into amortized initialization and per-online-step. Average of five trials. Preprocessing times amortized across $n$ operations.
$n$ Local (s) Client (s) Server (s) Client R. Total R.
1025 0.0751 0.232 0.121 3.08 4.69
2049 0.514 1.11 0.775 2.16 3.67
4097 4.46 5.79 6.15 1.30 2.68
8193 61.9 36.7 95.4 0.592 2.13
16385 511 200 675 0.392 1.71
Figure 2: Matrix-Matrix performance benchmarks. Average of five trials.
We see substantial performance improvements as $n$ grows in both contexts, but especially in matrix-vectors. Here, we achieve client improvements even at small $n$, with improvements increasing to factor-$25$ speedup for the client as we grow the matrix. Initialization costs are consistently negligible (due to the advantage of matrix-matrix over repeated matrix-vectors). In matrix-matrix, client improvements are less substantial, in large part due to the locality efficiencies available in dense (but not sparse) matrix multiplication. However, clients still achieve substantial advantage for medium-size matrices, up to nearly factor-three for $n=2^{14}$.
To the best of our knowledge, our method is the first to provide practical efficiency gains for the client in this parameter regime when encryption time is taken into account. Our efficiency is over two orders of magnitude higher than that of [BCH+24], despite encrypting both multiplicands while their approach encrypts one (note also that almost all of their cost is incurred by encryption/decryption). [^7] Their work, in turn, significantly outperforms [LZ22]. Our algorithm may be straightforwardly modified to only encrypt one multiplicand; this results in a significant additional improvement in client performance and server overhead.
Performance comparison to the later work of [BCH+25] is more complex. [^8] Our initial draft did not contain experimental data at their time of publication, so their work was unable to include experimental comparisons. We work over $\mathbb{Z}/2^{32}\mathbb{Z}$, while [BCH+25] work over $\mathbb{F}_{p}$ for $p$ a $32$-bit prime (and compares to plaintext operations over the same), rendering relative-to-plaintext performance not comparable. In addition, cryptographic assumptions need to be carefully evaluated to render accurate comparisons. Superficially, [BCH+25] obtain significant performance improvements for smaller values of $n$ ($\leq 2048$), while our work produces major gains for large $n$ ($\geq 4096$). In addition, [BCH+25] offers significant savings on client-side storage. The existence of constructions with tradeoffs suggests potential for future improvement.
We do not know of an efficient and well-motivated method for parameter search for optimal subspace dimensions/sparsities. Instead, we use a heuristic: at each step, aim for a constant ratio of subspace dimension to sparsity. We performed (fairly sparse) grid search for the optimal ratio for each $n$; therefore, our results merely upper-bound optimal runtimes.
9 Conclusion and Further Questions
We present natural definitions of Trapdoored-Matrix and Targeted-Trapdoored-Matrix objects, a method to turn instances of these objects into protocols for secure remote linear algebra, and instances of these objects that can be optimized to yield highly efficient versions of these protocols. It is not clear that these are the only (or best) primitives for this purpose, especially in the case of Trapdoored-Matrix, where evidence suggests the existence of notably faster pseudorandom matrices. Variants of our schemes may also be extensible to more general problems; harder problems in linear algebra, such as matrix polynomial evaluation, are a reasonable next step. Achieving a variety of operations parallel to that of [BCG+20a, CD23], or even [BCH+24] (while achieving practical efficiency gains) is a significantly more challenging problem, but would yield even broader applications.
{credits}
9.0.1 Acknowledgements
Research supported in part by the NSF Alan T. Waterman Award, Grant No. 1933331.
Thanks to Alex Lombardi for helpful discussion on LPN and related topics.
Appendix 0.A Performance Analysis of Algorithm 2
0.A.1 Asymptotic Analysis
We present a high-level exposition of performance here. An analysis of exact operation counts is available in Appendix 0.A.2.
Fix $\delta,\epsilon,\lambda$. Let $\nu$ be as stated in Section 3. We assume $m,n\geq\nu$. Set $n_{i}=\delta n_{i-1}$ for some $\delta<\frac{1}{2}$, $\mu_{i}=n_{i-1}^{\epsilon-1}$, $n_{d}=\nu$. We have $\sum_{i=1}^{d}n_{i}\leq\frac{\delta n}{1-\delta}$, $\mu_{i}n_{i-1}\leq n_{i-1}^{\epsilon}$. We set $\lambda^{\prime}=\frac{\lambda}{\lg|R|}$ to achieve statistical $\lambda$-bit security in our Mat-Mul-Check calls.
To clarify the following expressions, we omit the costs of the many calls to Mat-Mul-Check in the intermediate expressions, as they are never dominant.
The preprocessing runtime is dominated by the computation of the products $L_{1}\dots L_{i}$, the Right-Hidden-Mult call (which dominates the init call), and the construction of $A^{\prime}$. The cost of the first is none (apart from the above check cost) to the client and $O(n^{\omega}\delta^{2\omega-3})$ to the server. The cost of the second is $O(\nu^{\omega-2}(m+n)(\delta n)+(m+n)n^{1+\epsilon})$ to the client and $O\left(n^{\omega}\delta^{\omega-1}+\min(m,\delta n)^{\omega-2}n\max(m,\delta n)\right)$ to the server. The construction of $A$ costs $O(mn\nu+mn^{1+\epsilon})$ to the client. [^9] In many expressions, $\nu$ can often be tightened to $\nu^{\omega-2}$ via fast matrix multiplication, but this is uninteresting and complicates the resulting expressions. In total, the client cost is $O\left(n\left(n^{\epsilon}(m+n)+\nu(\delta n+m)\right)\right)$ and the server cost is $O\left(n^{\omega}\delta^{2\omega-3}+\min(m,\delta n)^{\omega-2}n\max(m,\delta n)\right)$.
The online step runtime for the client is $O((m+n)l\nu+(m+n)n^{\epsilon}l)$ to construct $B^{\prime}$, $AB^{\prime}$, and $A^{\prime}(B+B^{\prime})$, and $O(ml)$ for the final subtractions. The server performs one $m\times n$-by-$n\times l$ multiplication in the final line and incurs additional cost $O(n\max(\delta n,l)\min(\delta n,l)^{\omega-2})$.
We expound on two special cases:
0.A.1.1 Matrix Multiplication
One natural case is that with $m=n=l$ and one online step (i.e. a once-off matrix-matrix multiplication). Here, total client cost is $O\left(n^{2}\left(n^{\epsilon}+\nu\right)\right)$ and total server cost is, apart from one $n\times n$-by-$n\times n$ matrix multiplication, $O(n^{\omega}\delta^{\omega-2})$. In particular, our algorithm is $\left(n^{\epsilon}+\frac{\nu}{n},1+O(\delta^{\omega-2})\right)$-remote-efficient, resulting in a multiplicative $n^{1-\epsilon}$ saving for the client (ignoring the additive term in security parameter) and small overhead for the server.
0.A.1.2 Online Matrix-Vectors
Another natural case is $m=n,l=1$, and $k$ online steps. Here, setup cost is the same $O\left(n^{2}\left(n^{\epsilon}+\nu\right)\right)$ for the client and $O\left(\delta n^{\omega}\right)$ for the server, which amortize across $n^{\epsilon}$ and $\delta n$ online rounds, respectively. The per-multiplication online cost is $O(n(n^{\epsilon}+\nu))$ for the client and, apart from one $n\times n$ matrix-vector multiplication, $O(\delta n^{2})$ for the server. We therefore obtain (after amortization) the same $\left(n^{\epsilon}+\frac{\nu}{n},1+O(\delta^{\omega-2})\right)$-remote-efficiency guarantee and same small overheads. Note that the amortization occurs in a sublinear number of online steps; this holds because the initial multiplication benefits from fast matrix multiplication, while the online steps do not.
0.A.2 Per-Operation Analysis
Let $\mathfrak{a}(x)$ be the cost of adding/subtracting two $x$-element tensors; $\tilde{\mathfrak{d}}(x)$ be the cost of taking the dot product of two vectors, one of which is $x$-sparse; and $\mathfrak{m}(a,b,c)$ be the cost of multiplying two matrices, one of which is $a\times b$ and the other of which is $b\times c$. All operations are relative to our working ring.
We analyze the performance componentwise and with precise constants, as this algorithm exhibits particularly low constants. For simplicity, we omit summation terms, as these are negligible and/or dominated by adjacent dot-like operations, and random generation components, as these are fast in theory and practice and the total quantity of random entries generated is
$O\left(\sum_{i=1}^{d}nn_{i}+(m+l)n\mu_{1}+(m+l)n_{d}\right).$
Mat-Mul-Check:
Costs exactly $\mathfrak{m}(b,c,\lambda^{\prime})+\mathfrak{m}(a,b,\lambda^{\prime})+\mathfrak{m}(a,c,\lambda^{\prime})$ (plus vector equality checks, which are negligible in comparison).
Get-Checked-Cum-Prods:
By the above, costs
$\sum_{i=1}^{k-1}\mathfrak{m}(a_{i+1},a_{i+2},\lambda^{\prime})+\mathfrak{m}(a_{i},a_{i+1},\lambda^{\prime})+\mathfrak{m}(a_{i},a_{i+2},\lambda^{\prime})$
for the client and $\sum_{i=1}^{k}\mathfrak{m}(a_{1},a_{i+1},a_{i+2})$ for the server. Note that the first index in both sums is dominant in all uses.
Right-Hidden-Init:
By the above, costs $\sum_{i=0}^{d-1}\mathfrak{m}(s,n_{i},n_{i+1})$ for the server and $\mathfrak{m}(n_{0},n_{1})+\mathfrak{m}(s,n_{0})+\mathfrak{m}(s,n_{1})+\sum_{i=0}^{k-2}\mathfrak{m}(n_{i+1},n_{i+2},\lambda^{\prime})+\mathfrak{m}(n_{i},n_{i+1},\lambda^{\prime})+\mathfrak{m}(n_{i},n_{i+2},\lambda^{\prime})$ for the client.
Right-Hidden-Mult:
Computation of $V^{\prime}$ costs $\mathfrak{m}{n,n_{d},o}+\sum_{i=1}^{d}no\tilde{\mathfrak{d}}(n_{i-1}\mu_{i})$, and computation of $MV^{\prime}$ costs $\mathfrak{m}{s,n_{d},o}+\sum_{i=1}^{d}so\tilde{\mathfrak{d}}(n_{i-1}\mu_{i})$. The product call costs $\mathfrak{m}(s,n,o)$ for the server and $\mathfrak{m}(s,n,\lambda^{\prime})+\mathfrak{m}(n,o,\lambda^{\prime})+\mathfrak{m}(s,o,\lambda^{\prime})$ for the client.
Hidden-Init:
Calculating the cumulative products costs $\sum_{i=0}^{k-2}\mathfrak{m}(n_{i+1},n_{i+2},\lambda^{\prime})+\mathfrak{m}(n_{i},n_{i+1},\lambda^{\prime})+\mathfrak{m}(n_{i},n_{i+2},\lambda^{\prime})$ for the client and $\sum_{i=1}^{k}\mathfrak{m}(n_{1},n_{i+1},n_{i+2})$ for the server. The right-hidden steps are invoked with $s=\sum_{i=1}^{d}n_{i},o=m$. Computation of $A^{\prime}$ costs $\mathfrak{m}(m,n_{d},n)+\sum_{i=0}^{d-1}mn\tilde{\mathfrak{d}}(\mu_{i}n_{i-1})$ for the client.
Hidden-Mult:
Joint computation of $B^{\prime}$ and $AB^{\prime}$ costs $\mathfrak{m}(l,n_{d},n+m)+\sum_{i=0}^{d-1}l(n+m)\tilde{\mathfrak{d}}(\mu_{i}n_{i-1})$. The server pays $\mathfrak{m}(\sum_{i=1}^{d}n_{i},n,l)$ for the in-the-clear computation of the $L$-$(B+B^{\prime})$ products. The client then pays $\mathfrak{m}(m,n_{d},l)+\sum_{i=0}^{d-1}ml\tilde{\mathfrak{d}}(\mu_{i}n_{i-1})$. In the final line, the server pays $\mathfrak{m}(m,n,l)$ and the client pays $2\mathfrak{a}(ml)$.
Appendix 0.B Benchmarks for power-of-two $n$
$n$ Local (s) Client Pre. (s) Server Pre. (s) Client (s) Server (s) Client R. Total R.
1024 0.00523 0.0000746 0.0000315 0.00396 0.00472 0.773 1.68
2048 0.0239 0.000155 0.0000743 0.0121 0.0238 0.513 1.51
4096 0.104 0.000420 0.000402 0.0138 0.106 0.137 1.16
8192 0.422 0.000932 0.00172 0.0300 0.429 0.0732 1.09
16384 1.65 0.00279 0.00849 0.0826 1.66 0.0519 1.06
Figure 3: Variant of Table 1 without offset $n$. Note that some server time (operand transfer between client in server) is attributed to the client, hence server times slightly below naive time for small operand sizes.
$n$ Local (s) Client (s) Server (s) Client R. Total R.
1024 0.0716 0.239 0.112 3.33 4.90
2048 0.467 1.10 0.720 2.35 3.89
4096 5.22 6.03 5.96 1.15 2.30
8192 60.0 35.5 88.0 0.592 2.06
16384 503 190 812 0.378 1.99
Figure 4: Variant of Table 2 without offset $n$.